Year 2012 2013 2014 Adult Population 2,000 3,000 3,200 Number of Employed 1,400 1,300 1,600 Number...

Year

2012

2013

2014

Adult Population

2,000

3,000

3,200

Number of Employed

1,400

1,300

1,600

Number of Unemployed

200

600

200

The number of adults not in the labor force of Aridia in 2012 was

Group of answer choices

1,800

400

600

200
Answer #1

Answer) Labor force is sum of employed and unemployed(seeking job actively) adult population,in 2012 adult population was 2,000;out of which employed were 1400 and unemployed were 200,so total labor force becomes=Employed+Unemployed=1400+200=1600,so number of adults not in labor force=Adult population-labor force=2000-1600=400.

So correct answer is option b) 400

Answer is complete.Thank you!